OpenStudy (anonymous):

can you help me please? 7/10m - 17 = -1/2m -5

OpenStudy (anonymous):

\[\frac{ 7 }{ 10 }m -17 = -\frac{ 1 }{ 2 }m - 5\]

OpenStudy (anonymous):

solve for m?

OpenStudy (anonymous):

yah i know that but i dont know how to do that with a fraction can you show me how?

OpenStudy (anonymous):

@BlankDesign

OpenStudy (anonymous):

do you guys know

OpenStudy (anonymous):

it is multiple choice

OpenStudy (anonymous):

im sorry i dont know how

OpenStudy (anonymous):

M = 10 \[\frac{ 7 }{ 10 }m-17+\frac{ 1 }{ 2 }m=-5\] add 17 to both sides \[\frac{ 7 }{ 10 }m+\frac{ 1 }{ 2 }m=12\] make the fractions the same \[\frac{ 7 }{ 10 }m+\frac{ 5 }{ 10 }m=12\] add them together \[\frac{ 12 }{ 10 }m=12\] times 10 12m=120 divide by 12 m=10

OpenStudy (anonymous):

i know that you dont know how your out in thee living room!!!

OpenStudy (anonymous):

@DanL33 your amazing thanks for the help

OpenStudy (anonymous):

@ashlynn143 you're welcome